Two closed pipe produces 10 beats per second when emitting their fundamental nodes. If their lengths are in the ratio of 25: 26 their fundamental frequency in Hz are
(a) 270, 280
(b) 260, 270
(c) 260, 250
(d) 240, 250

Grade:Upto college level

2 Answers

Arun
3 years ago
We know that the fundamental frequency of a closed end pipe is given by
 
f=4Lv
 
f1=4L1v and f2=4L2v
 
f2f1=L1L2=2526
 
f1=2526f2
 
Now, beat=f1f2=10
 
2526f2f2=10
 
f2=250Hz
 
And, f1=2526f2=260Hz
 
Answer-(C)
Vikas TU
3 years ago
Dear student 
Above ans is not understandable 
f = v/4L 
f1 = v/4L1 
f2 = v/4L2 
f1/f2 = 26/25 
f1 - f2 =10 
26/25 f2 - f2 = 10 
f2 = 250 Hz 
f1 = 26/25 f2 = 260 Hz